OpenStudy (anonymous):

the ratio of the length and the width of the floor of a building is 7:4. they differ by 45 ft. what's the floor area of the building?

hartnn (hartnn):

let the length be 7x so width = 4x difference = 45 so 7x-4x=45 can u find x from here?

OpenStudy (anonymous):

let l = w+45 thus l:w= w+45:w =7:4 Can you simplify this?

OpenStudy (anonymous):

yep. thanks!

hartnn (hartnn):

so what area did u get ?

OpenStudy (anonymous):

.